A random sample of n1 = 16 communities in western Kansas gave the following information for people under 25 years of age.

x1: Rate of hay fever per 1000 population for people under 25

100 91 122 127 92 123 112 93
125 95 125 117 97 122 127 88

A random sample of n2 = 14 regions in western Kansas gave the following information for people over 50 years old.

x2: Rate of hay fever per 1000 population for people over 50

93 109 100 97 111 88 110
79 115 100 89 114 85 96

(i) Use a calculator to calculate x1, s1, x2, and s2. (Round your answers to two decimal places.)

x1 =
s1 =
x2 =
s2 =


(ii) Assume that the hay fever rate in each age group has an approximately normal distribution. Do the data indicate that the age group over 50 has a lower rate of hay fever? Use α = 0.05.
(a) What is the level of significance?


State the null and alternate hypotheses.

H0: μ1 = μ2; H1: μ1 < μ2H0: μ1 = μ2; H1: μ1μ2    H0: μ1 = μ2; H1: μ1 > μ2H0: μ1 > μ2; H1: μ1 = μ2


(b) What sampling distribution will you use? What assumptions are you making?

The Student's t. We assume that both population distributions are approximately normal with unknown standard deviations.The standard normal. We assume that both population distributions are approximately normal with known standard deviations.    The Student's t. We assume that both population distributions are approximately normal with known standard deviations.The standard normal. We assume that both population distributions are approximately normal with unknown standard deviations.


What is the value of the sample test statistic? (Test the difference μ1μ2. Round your answer to three decimal places.)


(c) Find (or estimate) the P-value.

P-value > 0.2500.125 < P-value < 0.250    0.050 < P-value < 0.1250.025 < P-value < 0.0500.005 < P-value < 0.025P-value < 0.005
Answer #1

a)

i)

x̅1=   109.750
s1 =    15.25
x̅2=   99.000
s2 =    11.48

ii)

Level of Significance ,    α =    0.05
H0: μ1 = μ2; H1: μ1 > μ2

b) The Student's t. We assume that both population distributions are approximately normal with unknown standard deviations

difference in sample means =    x̅1-x̅2 =    109.7500   -   99.0   =   10.750  
                          
pooled std dev , Sp=   √([(n1 - 1)s1² + (n2 - 1)s2²]/(n1+n2-2)) =    13.6316                  
std error , SE =    Sp*√(1/n1+1/n2) =    4.9887                  
                          
t-statistic = ((x̅1-x̅2)-µd)/SE = (   10.7500   -   0   ) /    4.99   =   2.155

c)

Degree of freedom, DF=   n1+n2-2 =    28

0.005 < P-value < 0.025